OpenStudy (anonymous):

complete the square to find vertex y=-x^2+5x−3

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

\[y= -x ^{2}+5x-3=-\left( x ^{2}-5x \right)-3\] \[y=-\left( x ^{2} -5x+ \left( \frac{ 5 }{ 2 } \right)^{2}-\left( \frac{ 5 }{2} \right)^{2}\right)-3\] \[y=-\left( x-\frac{ 5 }{ 2 } \right)^{2}+\frac{ 25 }{4 }-3\] \[y=-\left( x-\frac{ 5 }{ 2} \right)^{2}+\frac{ 13 }{4 }\]

OpenStudy (anonymous):

this is really hard to understand

OpenStudy (anonymous):

but Im still kind of confused :(

OpenStudy (anonymous):

keep -3 outside bracket

OpenStudy (anonymous):

then??

OpenStudy (anonymous):

the vertex??

OpenStudy (anonymous):

(-5/2, 37/4)

OpenStudy (anonymous):

thats the vertex right??

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

ok please re-state it

OpenStudy (anonymous):

just tell me the vertex !!!

OpenStudy (anonymous):

the vertex is ??? what

OpenStudy (anonymous):

-5/2, 37/4

OpenStudy (anonymous):

and??

OpenStudy (anonymous):

thanks!!

OpenStudy (anonymous):

it is not \[x ^{2}+5x-3\] but it is \[-x ^{2}+5x-3\]

OpenStudy (anonymous):

it is a downward parabola. in this case vertex is \[\left( \frac{ 5 }{ 2} ,\frac{ 13 }{ 4 }\right)\]